Let R be the region in the first quadrant bounded by the lines y= -2x+4,y= -2x+7,y=x-2 and y=x+1.By making the substitutions u=2x+y,v=x-y and integrating over a suitable region in the u-v plane. Evaluate the integral ∫∫(2x2-xy-y2)dxdy


1
Expert's answer
2021-07-21T16:56:45-0400


The sides of parallelogram are "2x+y=4,\ \ 2x+y=7,\ \ x-y=2,\ \ x-y=-1" .

By making the substitutions "u=2x+y" , "v=x-y" , we get the limits on the integral "4\leq u\leq 7" and "-1\leq v\leq 2" .

Now, solve for "x" and "y" :

"u+v=(2x+y)+(x-y)=3x" , "\ \ x=\frac{u+v}{3}"

"u-2v=(2x+y)-(2x-y)=3y" , "\ y=\frac{u-2v}{3}" .

Compute the Jacobian:

"J(u,v)=\begin{vmatrix}
 \frac{\partial x}{\partial u} & \frac{\partial x}{\partial v} \\
 \frac{\partial y}{\partial u} & \frac{\partial y}{\partial v}
\end{vmatrix}=\begin{vmatrix}
 \tfrac{1}{3} & \tfrac{1}{3} \\
 \tfrac{1}{3} & -\tfrac{2}{3}
\end{vmatrix}=-\frac{2}{9}-\frac{1}{9}=-\frac{1}{3}" and "\big|J(u,v)\big|=\frac{1}{3}"

The original integrand becomes "2x^2-xy-y^2=\frac{2(u+v)^2-(u+v)(u-2v)-(u-2v)^2
}{9}=uv" .

The double integral is "\iint \limits_R(2x^2-xy-y^2)dx\ dy=\int\limits _{-1}^{2}\int\limits_{4}^7uv \cdot \frac{1}{3}\ du \ dv =\frac{1}{3} \int\limits _{-1}^{2}v dv \cdot \int\limits_{4}^7u\ du =\frac{1}{3}\cdot \frac{v^2}{2}\bigg|^2_{-1} \cdot \frac{u^2}{2}\bigg|_4^7=\frac{1}{3}\cdot \frac{3}{2}\cdot \frac{33}{2}=\frac{33}{4}=8.25"


Answer: "\iint \limits_R(2x^2-xy-y^2)dx\ dy=8.25"